# Convergence Analysis and Acceleration of Fictitious Play for General Mean-Field Games via the Best Response

### v1

@@ 1 Introduction
The mean-field game (MFG) studies the Nash Equilibrium of a non-cooperative game involving a continuum of players. Since Larsy and Lions [37] and Huang, Malhamé and Caines [33] introduced the model independently, researchers from various fields studied the application of MFGs in, for example, economics [4], epidemics [42], crowd motion [5, 22], data science [49, 29, 20], etc. Recently, people in the machine learning community have been interested in connections between MFG and generative models [52, 43, 32], and between MFG and reinforcement learning [50, 29, 38].
In a MFG, each player seeks to optimize a strategy that minimizes their individual cost, in response to a given state distribution of the entire player population. And the individual best strategies collectively shape a new state distribution of the player population. The equilibrium of a MFG is the fixed point of this iterative process. More precisely, at the Nash Equilibrium, the strategy that optimizes each player’s cost also generates the equilibrium state distribution. This equilibrium can be characterized by a set of partial differential equations, comprising a Hamilton-Jacobi-Bellman (HJB) equation for determining each player’s optimal control, alongside a Fokker-Planck (FP) equation that depicts the players’ state distribution. Specifically, the solution of the MFG system is the pair of $(\rho,\phi)$ solves
$-\partial_{t}\phi-\nu\Delta\phi+H(x,\nabla\phi(x,t))=f(x,\rho(t)),$ $\quad\phi(x,T)=g(x,\rho(T)),$ (1a)
@@
This paper focuses on an approach known as fictitious play [19]. This strategy decouples the system and facilitates us to solve the HJB and FP alternatively by simple time-marching methods. It has a natural interpretation in game theory and economics. In each iteration, the players seek the best strategy in response to the current state distribution. They then update the state distribution by taking a weighted average of the current distribution and the distribution resulting from the best strategy, which guides the next strategy search.
@@
An important study of fictitious play is its convergence analysis. For potential mean-field games (MFGs), the convergence has been established with the help of the potential functional, as shown in [19], and the sublinear and linear convergence rates for different $\delta_{k}$ weights were recently proven in [41]. However, for non-potential MFGs, since a variational structure is absent, the convergence analysis is more challenging and less understood.
@@
In this paper, we establish the convergence rate of fictitious play for general mean-field games (MFGs) through an analysis of the best-response mapping. First, we reveal an interesting geometric relation between the current state distribution and best response distribution it leads to: they lies on the ”opposite side” with respect to the equilibrium state in terms of the monotonicity of the interaction cost (see Lemmas 2.10 and 2.13), which reveals why a convex combination of the two provides a stable and closer approximation to the equilibrium in a fictitious play strategy. Next, we define the gain of the best response as the difference between the cost of the given distribution and the minimal cost achieved by the best response. By definition, this gain is always non-negative and reaches zero only when the given distribution is the Nash Equilibrium distribution. The gain of the best response, also known as exploitability in [30] and other works, serves as a key metric in our analysis. Through a stability study of the best-response mapping and its fixed point, we demonstrate that the gain of the best response is equivalent to the distance between the given distribution and the Nash Equilibrium. We further show that the gain of the best response acts as a Lyapunov function and decays to zero under the directional convexity of the dynamic cost and monotonicity of the interaction cost. It therefore indicates the convergence and provides a convergence rate of fictitious play for general MFGs. To the best of our knowledge, this is the first convergence rate established for fictitious play without assuming the existence of a potential.
@@
To accelerate convergence in fictitious play, we introduce a backtracking line search method to dynamically select the largest feasible weight $\delta_{k}$ at each iteration. Our convergence analysis shows that the convergence rate is explicitly influenced by the choice of $\delta_{k}$. While traditional fictitious play algorithms often use a predefined decay rate, such as $\delta_{k}=\frac{1}{k}$, this approach results in slower convergence as $k$ increases. In contrast, our backtracking line search approach adapts $\delta_{k}$ to maximize efficiency at each step. While previous work has applied similar line search techniques for potential MFGs using potential costs [41], our method is independent of potential structures and is therefore suitable for general MFGs.
@@
    We develop the best-response mapping and analyze its stability, which in turn establishes the stability of its fixed point, the Nash Equilibrium. This analysis yields a Lyapunov function that provides an explicit convergence rate for the fictitious play algorithm for a general class of mean-field games (MFGs). To the best of our knowledge, this is the first convergence rate established for non-potential MFGs.
@@
Different ways to solve for the optimal control and to update the density distribution lead to different implementations of fictitious play. In this paper, we focus on alternatively solving HJB and FP equations. Among the existing literature, the work presented in [34] closely aligns with our investigation. This study utilizes the Hopf-Cole transformation technique to linearize the MFG system, which was initially introduced in [16]. It then employs fictitious play combined with the finite difference method to tackle the linearized system. While [16] focuses on systems with quadratic Hamiltonians, [34] extends the methodology to MFGs with generalized quadratic Hamiltonians. In contrast, our paper adopts a direct discretization of the original MFG system, iterating between solving the HJB and FP equations. This approach enables us to work with general Hamiltonians.
When the state space is finite [26, 24], many works adopt reinforcement learning strategies to obtain the optimal control instead of solving the HJB equation. While this reinforcement learning can be inaccurate, [23] studies the error propagation in fictitious play and shows that fictitious play is still effective even when the model is not known and there are accumulated errors in the learning process. For more studies using reinforcement learning in fictitious play, we refer readers to the review paper [40].
Besides fictitious play, when the Hamiltonian is quadratic, [28] applies the Hopf-Cole transformation to linearize the system and then implements fixed-point iteration. They also prove that this adjusted process ensures convergence. For the general Hamiltonian, another strategy adapts the policy iteration in control theory to the MFG setting. It solves the HJB equation in two phases: initially freezing the action to linearize the HJB for solving the value function, followed by action adjustment based on this value function. [14, 15, 17] establish the convergence rate of the algorithm for polynomial growth Hamiltonian and [39] for non-separable Hamiltonian. [47] proposes and proves the convergence of smoothed policy iterations which use an average of policy to update the density in policy iteration.
@@
Multi-grid is an effective preconditioner that can accelerate an iterative solver for elliptic equations. The key idea is that the iterative solver is effective in reducing errors in high frequencies (relative to the grid size). So one can use multiple grid sizes to deal with all frequencies effectively. Some, but not extensive, literature has explored the application of the multi-grid idea to solve a MFG problem or its subproblem. [7] explores accelerating Newton’s methods for the coupled global nonlinear discretized system in space and time with multigrid and preconditioning. Additionally, [13, 8] apply the multigrid methods to solve the (bi-)Laplacian equation that arises in the subproblem of optimization algorithms like primal-dual and ADMM. All the above use of multi-grid is aimed as a preconditioner to deal with problems of elliptic nature. However, as noted in [7, 6], for second-order MFGs, the elliptic operator is second-order in time and fourth-order in space, leading to anisotropy that renders a full-coarsening multigrid approach ineffective. Therefore [7, 13, 8] only apply multigrid on the spatial domain for second-order MFGs. Our hierarchical grid strategy is totally different. The key idea of using a hierarchical grid is to accelerate information propagation from the boundary for a MFG as a two-point boundary value problem. Hence our strategy applies to both first-order and second-order MFGs. In addition, while multigrid deals with frequencies from high to low and hence goes from fine to coarse grid, our strategy always goes from coarse to fine, using the approximate solution on a coarse grid, which can achieve a coarse balance between boundary conditions at both ends more quickly and stably, as an initial guess to a finer grid to find a more accurate solution. Moreover, our hierarchical strategy automatically incorporates a numerical viscosity which is proportional to the grid size in our finite difference scheme. It allows a natural computation of the vanishing viscosity solution for the best response for first order MFGs which enhances the numerical stability. In this sense, the spirit of the multilevel approach in [51] is the closest to ours, where the interpolation of the solution on the coarse grid serves as a good initial guess on the fine grid. However, their method is an optimization approach for the global (in space and time) and coupled discrete system based on the variational formulation. It can not be applied to non-potential MFGs. Also, due to the use of straightforward numerical discretization for the HJB equation, the stability can be a problem for first-order MFGs or MFGs with small diffusion. To the best of our knowledge, our work is the first one employing the full-coarsening hierarchical grid and combining it with the decoupling approach.
@@
Many existing numerical methods for MFG view the system for both PDEs of $\phi$ and $\rho$ in space and time as a whole system and aim to solve it altogether. For example, [3] applies finite difference to the MFG system and solves the entire system with Newton solver. A more common approach focuses on potential MFG with a variational formulation and applies an optimization solver to it, see [9, 10, 44, 51] for low-dimensional cases and [43, 46, 32] for high-dimensional cases. While this coupled approach directly handles the global structure of the PDE system, the full discretization in space and time results in a large nonlinear system to solve. Moreover, as the grid size becomes finer, the nonlinear system not only becomes larger but also worse conditioned due to the differential operation. In general, a direct solver ignoring the physics and structures of the underlying PDE system can be inefficient. Besides computational concerns, all optimization strategies rely on the equivalent variational formulation of the problem and are not easy to adapt to non-potential MFG problems.
@@
The rest of this paper is organized as follows. Section 2 begins by outlining our problem setting and reviewing the formulation of mean-field games (MFGs) and the fictitious play algorithm. It then provides a detailed convergence analysis for both potential and non-potential MFGs. The details of our discretization and implementation are in Section 3. Section 4 verifies the convergence analysis and demonstrates the efficacy of our proposed schemes through numerical experiments. Finally, Section 5 concludes the paper and discusses future works. Our main contributions are the convergence analysis in Section 2.2, especially Sections 2.2.2 and 2.2.3, and the accelerations in Section 2.2.4 and 3.2.
2 Fictitious Play for Mean-Field Games
In this section, we first review MFGs and the fictitious play algorithm and then provide the convergence analysis and the back-tracking line search acceleration of fictitious play. We slightly abuse our notation and use $\rho$ to represent a probability distribution or its density function, depending on the context.
2.1 A brief review of MFG and fictitious play

### v2

@@ 1 Introduction
The mean-field game (MFG) studies the Nash Equilibrium of a non-cooperative game involving a continuum of players. Since Larsy and Lions lasry2007mfg and Huang, Malhamé and Caines huang2006mfg introduced the model independently, researchers from various fields studied the application of MFGs in, for example, economics achdou2022income , epidemics lee2021controlling , crowd motion achdou2019mean ; degond2014large , data science weinan2019mean ; guo2019learning ; carmona2023model , etc. Recently, people in the machine learning community have been interested in connections between MFG and generative models zhang2023meanfieldgenerative ; Lin2021apacnet ; huang2023nfmfg , and between MFG and reinforcement learning yang2018mean ; guo2019learning ; lauriere2022scalable .
In an MFG, each player seeks to optimize a strategy that minimizes their individual cost, in response to a given state distribution of the entire player population. The individual best strategies collectively shape a new state distribution of the player population. The equilibrium of an MFG is the fixed point of this iterative process. More precisely, at the Nash Equilibrium, the strategy that optimizes each player’s cost also generates the equilibrium state distribution. This equilibrium can be characterized by a set of partial differential equations, comprising a Hamilton-Jacobi-Bellman (HJB) equation for determining each player’s optimal control, alongside a Fokker-Planck (FP) equation that depicts the players’ state distribution. Specifically, the solution of the MFG system is the pair of $(\rho,\phi)$ that solves
$-\partial_{t}\phi-\nu\Delta\phi+H(x,\nabla\phi(x,t))=f(x,\rho(t)),$ $\quad\phi(x,T)=f_{T}(x,\rho(T)),$ (1a)
@@
This paper focuses on an approach known as fictitious play cardaliaguet2017ficplay . This strategy decouples the system and facilitates us to solve the HJB and FP alternatively by simple time-marching methods. It has a natural interpretation in game theory and economics. In each iteration, the players seek the best strategy in response to the current state distribution. They then update the state distribution by taking a weighted average of the current distribution and the distribution resulting from the best strategy, which guides the next strategy search.
@@
An important study of fictitious play is its convergence analysis. For potential mean-field games (MFGs), the convergence has been established with the help of the potential functional, as shown in cardaliaguet2017ficplay , and the sublinear and linear convergence rates for different $\delta_{k}$ weights were recently proven in Lavigne2023gcg-ficplay . However, for non-potential MFGs, since a variational structure is absent, the convergence analysis is more challenging and less understood.
@@
In this paper, we establish the convergence rate of fictitious play for general mean-field games (MFGs) through an analysis of the best-response mapping. First, we reveal an interesting geometric relation between the current state distribution and best response distribution it leads to: they lie on the ”opposite side” with respect to the equilibrium state, assuming the monotonicity of the interaction cost (see Proposition 3.11). This observation explains why direct fixed-point iteration may fail and why fictitious play yields a more stable and accurate approximation to the equilibrium. Next, we study the stability of the best response. Although the best-response objective is typically convex but not uniformly strongly convex, we show that it exhibits directional strong convexity between any two distinct function pairs in the feasible set (see Section 3.2). This observation ensures the stability of the best response mapping. We then define the gain of the best response as the difference between the cost of the given distribution and the minimal cost achieved by the best response. This is also known as exploitability in hadikhanloo2017anonymous . By definition, the gain of the best response is always non-negative and vanishes only when the given distribution corresponds to a Nash Equilibrium (see Lemma 3.4). Our main Theorem 3.19 shows that the gain of the best response acts as a Lyapunov function and decays to zero under the directional convexity of the dynamic cost and monotonicity of the interaction cost. It therefore indicates the convergence and provides a convergence rate of fictitious play for general MFGs. To the best of our knowledge, this is the first convergence rate established for fictitious play without assuming the existence of a potential.
@@
To accelerate convergence in fictitious play, we introduce a backtracking line search method to dynamically select the largest feasible weight $\delta_{k}$ at each iteration. Our convergence analysis shows that the convergence rate is explicitly influenced by the choice of $\delta_{k}$. While traditional fictitious play algorithms often use a predefined decay rate, such as $\delta_{k}=\frac{1}{k}$, this approach results in slower convergence as $k$ increases. In contrast, our backtracking line search approach adapts $\delta_{k}$ to maximize efficiency at each step. While previous work has applied similar line search techniques for potential MFGs using potential costs Lavigne2023gcg-ficplay , our method is independent of potential structures and is therefore suitable for general MFGs.
@@
    We connect the Nash Equilibrium and fictitious play with the best response mapping. The best response yields a Lyapunov function, and the stability analysis of the best response provides an explicit convergence rate for the fictitious play algorithm for a general class of mean-field games (MFGs). To the best of our knowledge, this is the first convergence rate established for non-potential MFGs.
@@
Different ways to solve for the optimal control and to update the density distribution lead to different implementations of fictitious play. In this paper, we focus on alternatively solving HJB and FP equations. Among the existing literature, the work presented in inoue2023ficplaylsmfg closely aligns with our investigation. This study utilizes the Hopf-Cole transformation technique to linearize the MFG system, which was initially introduced in gueant2012mean . It then employs fictitious play combined with the finite difference method to tackle the linearized system. While gueant2012mean focuses on systems with quadratic Hamiltonians, inoue2023ficplaylsmfg extends the methodology to MFGs with generalized quadratic Hamiltonians. In contrast, our paper adopts a direct discretization of the original MFG system, iterating between solving the HJB and FP equations. This approach enables us to work with general Hamiltonians.
When the state space is finite gomes2013continuous ; Gao2024finitemfg , many works adopt reinforcement learning strategies to obtain the optimal control instead of solving the HJB equation. While this reinforcement learning can be inaccurate, elie2020convergence studies the error propagation in fictitious play and shows that fictitious play is still effective even when the model is not known and there are accumulated errors in the learning process. For more studies using reinforcement learning in fictitious play, we refer readers to the review paper laurière2024learningreview .
Besides fictitious play, when the Hamiltonian is quadratic, gueant2012mean applies the Hopf-Cole transformation to linearize the system and then implements fixed-point iteration. They also prove that this reformulation of fixed-point iteration converges. For the general Hamiltonian, another strategy adapts the policy iteration in control theory to the MFG setting. It solves the HJB equation in two phases: initially freezing the action to linearize the HJB for solving the value function, followed by action adjustment based on this value function. cacase2021policyitercvg ; camilli2022policyitercvg ; camilli2022policyitercvgrate establish the convergence rate of the algorithm for polynomial growth Hamiltonian and Laurière2023nonseparableHam for non-separable Hamiltonian. tang2024smpolicyiter proposes and proves the convergence of smoothed policy iterations, which use an average of the policy to update the density in policy iteration.
@@
Multi-grid is an effective preconditioner that can accelerate an iterative solver for elliptic equations. The key idea is that the iterative solver is effective in reducing errors in high frequencies (relative to the grid size). So one can use multiple grid sizes to deal with all frequencies effectively. Some, but not extensive, literature has explored the application of the multi-grid idea to solve an MFG problem or its subproblem. achdou2012multigrid explores accelerating Newton’s methods for the coupled global nonlinear discretized system in space and time with multigrid and preconditioning. Additionally, briceno2018multigridadmm ; andreev2017multigridadmm apply the multigrid methods to solve the (bi-)Laplacian equation that arises in the subproblem of optimization algorithms like primal-dual and ADMM. All the above use of multi-grid is aimed as a preconditioner to deal with problems of elliptic nature. However, as noted in achdou2012multigrid ; AchdouLauriere2020review , for second-order MFGs, the elliptic operator is second-order in time and fourth-order in space, leading to anisotropy that renders a full-coarsening multigrid approach ineffective. Therefore achdou2012multigrid ; briceno2018multigridadmm ; andreev2017multigridadmm only apply multigrid on the spatial domain for second-order MFGs. Our hierarchical grid strategy is totally different. The key idea of using a hierarchical grid is to accelerate information propagation from the boundary for an MFG as a two-point boundary value problem. Hence our strategy applies to both first-order and second-order MFGs. In addition, while multigrid deals with frequencies from high to low and hence goes from fine to coarse grid, our strategy always goes from coarse to fine, using the approximate solution on a coarse grid, which can achieve a coarse balance between boundary conditions at both ends more quickly and stably, as an initial guess to a finer grid to find a more accurate solution. Moreover, our hierarchical strategy automatically incorporates a numerical viscosity that is proportional to the grid size in our finite difference scheme. It allows a natural computation of the vanishing viscosity solution for the best response for first-order MFGs which enhances the numerical stability. In this sense, the spirit of the multilevel approach in yu2024pgdeucmfg is the closest to ours, where the interpolation of the solution on the coarse grid serves as a good initial guess on the fine grid. However, their method is an optimization approach for the global (in space and time) and coupled discrete system based on the variational formulation. It can not be applied to non-potential MFGs. Also, due to the use of straightforward numerical discretization for the HJB equation, the stability can be a problem for first-order MFGs or MFGs with small diffusion. To the best of our knowledge, our work is the first one employing the full-coarsening hierarchical grid and combining it with the decoupling approach.
@@
Many existing numerical methods for MFG view the system for both PDEs of $\phi$ and $\rho$ in space and time as a whole system and aim to solve it altogether. For example, achdou2010mean applies finite difference to the MFG system and solves the entire system with Newton solver. A more common approach focuses on potential MFG with a variational formulation and applies an optimization solver to it, see benamou2015augmented ; benamou2017variational ; liu2021nonlocal ; yu2024pgdeucmfg for low-dimensional cases and Lin2021apacnet ; Ruthotto2020mfgmachinelearning ; huang2023nfmfg for high-dimensional cases. While this coupled approach directly handles the global structure of the PDE system, the full discretization in space and time results in a large nonlinear system to solve. Moreover, as the grid size becomes finer, the nonlinear system not only becomes larger but also worse conditioned due to the differential operation. In general, a direct solver ignoring the physics and structures of the underlying PDE system can be inefficient. Besides computational concerns, all optimization strategies rely on the equivalent variational formulation of the problem and are not easy to adapt to non-potential MFG problems.
@@
We summarize the notations in Table 1. The remainder of the paper is organized as follows. Section 2 introduces the problem setting and reviews the formulation of mean-field games (MFGs) along with the fictitious play algorithm. In Section 3, we reformulate the MFG and fictitious play dynamics using the concept of the best response, analyze the stability of the best response map, and provide a detailed convergence analysis for both potential and non-potential MFGs. Section 4 describes our discretization and implementation details. Section 5 verifies the convergence theory and demonstrates the effectiveness of the proposed methods through numerical experiments. Finally, Section 6 concludes the paper and discusses directions for future work.
Our main contributions lie in the convergence analysis developed in Section 3, with a focus on Sections 3.3 and 3.4, as well as in the proposed acceleration strategies presented in Sections 3.5 and 4.2. The analysis is built upon three key components: the Lyapunov function defined in (23), the qualitative property established in Proposition 3.11, and the notion of directional strong convexity in Section 3.2.
For simplicity, we focus on second-order MFGs ($\nu>0$) and periodic boundary conditions in Sections 2 and 3. We believe our results naturally extend to problems posed on $\mathbb{R}^{d}$ or on a bounded domain with suitable boundary conditions. Although first-order MFGs ($\nu=0$) involve technical subtleties, we expect our results to hold when the solution is sufficiently smooth and $\rho>0$ throughout. To illustrate this, in Section 4, we detail a discretization on a bounded domain with homogeneous Neumann conditions for the HJB equation and no-flux conditions for the Fokker–Planck equation, which are used in the experiments in Section 5.
Table 1: Summary of Notations
$\mathbb{T}^{d}=\mathbb{R}^{d}/\mathbb{Z}^{d}$ is the space domain, with $L^{2}$ inner product $\langle\cdot,\cdot\rangle_{\mathbb{T}^{d}}$ and $L^{2}$ norm $\|\cdot\|_{\mathbb{T}^{d}}$.
$T>0$ is the given terminal time.
$Q_{T}:=\mathbb{T}^{d}\times(0,T)$ is the joint domain, with $L^{2}$ inner product $\langle\cdot,\cdot\rangle_{Q_{T}}$ and $L^{2}$ norm $\|\cdot\|_{Q_{T}}$.
$\mathcal{P}(\mathbb{T}^{d})$ is the set of Borel probability measures on $\mathbb{T}^{d}$.
$\rho=(\rho_{t})_{t\in[0,T]},\rho_{t}\in\mathcal{P}(\mathbb{T}^{d})$ is the distribution/density evolution in most cases. Depending on context, $\rho\in[0,+\infty)$ denotes the density at one point, or $\rho\in\mathcal{P}(\mathbb{T}^{d})$ denotes a probability distribution or its density.
$v:\mathbb{T}^{d}\times[0,T]\to\mathbb{R}^{d}$ is the velocity field in most cases. Depending on context, $v\in\mathbb{R}^{d}$ is a vector at one point, or $v:\mathbb{T}^{d}\to\mathbb{R}^{d}$ is the velocity field at some time $t$.
$m:\mathbb{T}^{d}\times[0,T]\to\mathbb{R}^{d},m=\rho v$ is the momentum in most cases. Depending on context, $m\in\mathbb{R}^{d}$ is a vector at some $(x,t)$, or $m:\mathbb{T}^{d}\to\mathbb{R}^{d}$ is the velocity field at some time $t$.
$H:\mathbb{T}^{d}\times\mathbb{R}^{d}\to\mathbb{R},(x,p)\mapsto H(x,p)$ is the Hamiltonian.
$L:\mathbb{T}^{d}\times\mathbb{R}^{d},(x,v)\mapsto L(x,v)$ is the convex conjugate of $H$ defined in (7).
$f,f_{T}$ are interaction and terminal costs, either nonlocal $\mathbb{T}^{d}\times\mathcal{P}(\mathbb{T}^{d})\to\mathbb{R}$ or local $\mathbb{T}^{d}\times[0,+\infty)\to\mathbb{R}$.
$C$ is the constraint set of $(\rho,m)$, defined in (6).
$R:C\to\mathbb{R}$ is the dynamic cost, defined in (8).
$J(\cdot,\cdot;\rho):C\to\mathbb{R}$ is the optimal control cost under given density evolution $\rho$, defined in (16).
$g:C\to\mathbb{R}$ is the gain of the best response, defined in (23).
$F,F_{T}:\mathcal{P}(\mathbb{T}^{d})\to\mathbb{R}$ are the potential of $f,f_{T}$ if $f,f_{T}$ derive from potential.
$J:C\to\mathbb{R}$ is the potential cost of a potential MFG, defined in (10).
$h:C\to\mathbb{R}$ is the primal gap in potential MFG, defined in (86).
2 A Brief Review of MFG and Fictitious Play
Let the time domain be $[0,T]$ ($T>0$), the spatial domain be $\mathbb{T}^{d}$, and define the space-time domain as $Q_{T}:=\mathbb{T}^{d}\times(0,T)$. We denote the $L^{2}$ inner product and norm on $\mathbb{T}^{d}$ by $\langle\cdot,\cdot\rangle_{\mathbb{T}^{d}}$ and $\|\cdot\|_{\mathbb{T}^{d}}$, and those on $Q_{T}$ by $\langle\cdot,\cdot\rangle_{Q_{T}}$ and $\|\cdot\|_{Q_{T}}$. $\mathcal{P}(\mathbb{T}^{d})$ is the set of Borel probability measures on $\mathbb{T}^{d}$. The Hamiltonian $H$, interaction cost $f$, and terminal cost $g$ are periodic in space. The initial condition $\rho_{0}\in\mathcal{P}(\mathbb{T}^{d})$ is periodic in space. When $\rho\in\mathcal{P}(\mathbb{T}^{d})$ is absolutely continuous, we slightly abuse notation and use $\rho$ to refer to either a probability distribution or its density function, depending on the context. When $\rho$ refers to the density function, $\rho>0$ implies $\rho(x)>0$ for any $x\in\mathbb{T}^{d}$.
2.1 Second-order MFGs and monotonicity
A second-order Mean Field Game (MFG) is the following system of $(\rho,\phi)$ with $\nu>0$:
[…]